I think that this is the topic. I need to solve for x in this equation.
15-2x=5
I don't know how do solve for x and the steps that you do inorder to get the correct answer.

Answer by checkley75(3666)

15-2X=5
-2X=5-15
-2X=-10
X=-10/-2
X=5 ANSWER.
PROOF
15-2*5=5
15-10=5
5=5